log[9]729=v value of v

Answer by stanbon(75887)

log[9]729=v
---
9^v = 729
v = 3
---------------------
Another way:
v = log9(729) = log(729)/log(9) = 2.8627../0.9542..
= 3
===============
Cheers,
stan H.